Forrester Fashions has annual credit sales of 250,000 units with an average collection period of 70 days. The company has a per-unit variable cost of $20 and a per-unit sale price of $30.00. Bad debts currently are 5% of sales. The firm estimates that a proposed relaxation of credit standards would not affect its 70-day average collection period but would increase bad debts to 7.50% of sales. which would increase to 300,000 units per year. Forrester requires a 12% return on investments. Show all necessary calculations required to evaluate Forresters proposed relaxation of credit standards. (Note: Assume a 365-day year.) The additional profit contribution from an increase in sales is $ (Round to the nearest dollar.)
Answer #1

Answer:

Current (before proposed relaxation in credit terms):

Credit sales = Units * sales price = 250,000 * $30 = $7,500,000

Average receivables = $7,500,000 * 70 /365 = $1,438,356.16

Contribution per unit = $30 - $20 = $9

Bad debt = $7,500,000 * 5% = $375,000

After proposed relaxation in credit:

Credit sales = Units * sales price = 300,000 * $30 = $9,000,000

Average receivables = $9,000,000 * 70 /365 = $1,726,027.40

Bad debt = $9,000,000 * 7.5% = $675,000

Incremental costs and benefits:

Benefit:

Increase in contribution = Increase in sales units * contribution per unit = (300,000 - 250,000) * $10 = $500,000

Costs:

Increase in bad debt = $675,000 - $375,000 = $300,000

Cost of capital due increase in receivables = ($1,726,027.40 - $1,438,356.16) * 12% = $34,420.55

Additional profit contribution from an increase in sales is = $500,000 - $300,000 - $34,420.55 = $165,579.45

Additional profit contribution from an increase in sales is $ 165,579